# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ Abstract
Many applications, e.g., in shared mobility, require coordinating a large number of agents. Mean-field reinforcement learning addresses the resulting scalability challenge by optimizing the policy of a representative agent. In this paper, we address an important generalization where there exist global constraints on the distribution of agents (e.g., requiring capacity constraints or minimum coverage requirements to be met). We propose Safe-$\text{M}^{3}$-UCRL, the first model-based algorithm that attains safe policies even in the case of unknown transition dynamics. As a key ingredient, it uses epistemic uncertainty in the transition model within a log-barrier approach to ensure pessimistic constraints satisfaction with high probability. We showcase Safe-$\text{M}^{3}$-UCRL on the vehicle repositioning problem faced by many shared mobility operators and evaluate its performance through simulations built on Shenzhen taxi trajectory data. Our algorithm effectively meets the demand in critical areas while ensuring service accessibility in regions with low demand.
1 Introduction
Multi-agent reinforcement learning (MARL) is a rapidly growing field that seeks to understand and optimize the behavior of multiple agents interacting in a shared environment. MARL has a wide range of potential applications, including vehicle repositioning in shared mobility services (e.g., moving idle vehicles from low-demand to high-demand areas [35]), swarm robotics (e.g., operating a swarm of drones), and smart grids (e.g., operating a network of sensors in electric system [46]). The interactions between agents in these complex systems introduce several challenges, including non-stationarity, scalability, competing learning goals, and varying information structure.
Mean-Field Control (MFC) addresses the scalability and non-stationarity hurdles associated with MARL by exploiting the insight that many relevant MARL problems involve a large number of very similar agents. Instead of focusing on the individual agents and their interactions, MFC considers an asymptotically large population of identical agents and models them as a distribution on the state space. The solutions obtained by MFC are often sufficient for the finite-agent equivalent problem [29, 40] in spite of the introduced approximations. An example of such a system is a ride-hailing platform that serves on-demand trips with a fleet of vehicles. The platform needs to proactively reposition idle vehicles based on their current locations, the locations of the other vehicles in the fleet, and future demand patterns (see Figure 1) to maximize the number of fulfilled trips and minimize customer waiting times. Additionally, the platform may be obligated by external regulators to guarantee service accessibility across the entire service region. The problem quickly becomes intractable as the number of vehicles increases.
In this paper, we focus on learning the safe optimal policies in a setting where the underlying dynamics are unknown. In most real-world systems, the dynamics must be learned from the data obtained from repeated interactions with the environment. We assume that the cost of obtaining data from the environment is high and seek to design a model-based solution that efficiently uses the collected data. Existing works consider solving the MFC problem via model-free or
@@ 1 Introduction
Figure 1: An illustrative example of taxis’ spatial distribution (light-blue scatters), repositioning trips (blue arrows), and a taxi’s trajectory as a sequence of passenger trips (red arrows).
model-based methods without safety guarantees. However, the proposed Safe Model-Based Multi-Agent Mean-Field Upper-Confidence RL (Safe-$\text{M}^{3}$-UCRL) algorithm focuses on learning optimal policies for MFC while avoiding undesired or dangerous distributions of agents’ population.
@@
Contributions. Section 3 considers a novel mean-field setting with safety constraints and defines a novel comprehensive notion of global population-based safety. To address safety-constrained environments featuring a large population of agents, in Section 4, we propose a model-based mean-field reinforcement learning algorithm called Safe-$\text{M}^{3}$-UCRL. Our algorithm leverages epistemic uncertainty in the transition model, employing a log-barrier approach to guarantee pessimistic satisfaction of safety constraints and enables the derivation of safe policies. In Section 5, we conduct empirical testing on the vehicle repositioning problem, a challenge commonly faced by shared mobility operators. Our results demonstrate that the policies derived using Safe-$\text{M}^{3}$-UCRL successfully fulfill demand in critical demand hotspots while ensuring service accessibility in areas with lower demand.
2 Related Work
@@
Our notion of safety for the mean-field problem extends the frameworks of Mean-Field Games (MFG) and Mean-Field Control (MFC) [31, 32, 26, 25]. For a summary of the progress focusing on MFGs, see [33] and references therein. We focus on MFC in this work which assumes cooperative agents in contrast to MFGs which assume competition. [19, 41, 22, 23, 5] address the problem of solving MFCs with known dynamics, i.e., planning, while [52, 9, 4, 3, 8, 53, 10, 51] consider model-free Q-Learning and Policy Gradient methods in various settings. Closest to our approach, [43] introduces $\text{M}^{3}$-UCRL, a model-based, on-policy algorithm, which is more sample efficient than other proposed approaches. Similarly to [28, 12, 15] for model-based single-agent RL and [47] for model-based MARL, $\text{M}^{3}$-UCRL uses the epistemic uncertainty of the transition dynamics model to design optimistic policies that efficiently balance exploration and exploitation in the environment and maximize sample efficiency. This is also the setting of our interest. However, there is no consideration of safety in any of these methods.
In terms of safety, there are two main ways of handling it in RL [18]; assigning significantly lower rewards to unsafe states [38] and providing additional knowledge to the agents [49] or using the notion of controllability to explicitly avoid unsafe policies [20]. Furthermore, the following approaches combine the two methods; [6] uses Lyapunov functions to restrict the safe policy space, [11] projects unsafe policies to a safe set via a control barrier function, and [2] introduces shielding, i.e., correcting actions only if they lead to unsafe states. The aforementioned works consider only individual agents, while in large-scale multi-agent environments, maintaining individual safety becomes intractable, and the focus shifts towards global safety measures. [57] show how to use the notion of returnability to define a safe, decentralized, multi-agent version of Q-learning, however, their approach does not scale well for large systems. [40] introduces constraints to the MFC by defining a cost function and a threshold that the discounted sum of costs can not exceed. We propose a different formulation that restricts the set of feasible mean-field distributions at every step, therefore, allowing for more specific control over constraints and safe population distributions.
3 Problem Statement
We consider the Mean-Field Control (MFC) model to capture a collective behavior of a large number of collaborative agents operating in the shared stochastic environment. This model assumes the limiting regime of infinitely many agents and homogeneity. Namely, all agents are identical and indistinguishable. Solving MFC amounts to finding an optimal policy for a single, so-called, representative agent due to the identical agent assumption. The representative agent interacts with the distribution of agents instead of focusing on individual interactions and optimizes a collective reward. Its policy is used to control all the agents in the environment.
@@ 3 Problem Statement
Formally, we consider the episodic setting, where episodes $n=1,\ldots,N$ each have $t=0,\ldots,T-1$ discrete steps. The state space $\mathcal{S}\subseteq\mathbb{R}^{p}$ and action space $\mathcal{A}\subseteq\mathbb{R}^{q}$ are the same for every agent. We use $s_{n,t}^{(i)}\in\mathcal{S}$ and $a_{n,t}^{(i)}\in\mathcal{A}$, to denote the state and action of agent $i\in\{1,\ldots,m\}$ in episode $n$ at step $t$. At every $n$ and $t$, the mean-field distribution $\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ describes the global state with $m$ identical agents when $m\rightarrow+\infty$:
$\mu_{n,t}(s)=\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{n,t}^{(i)}=s),$
@@
We posit that the reward $r:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathbb{R}$ of the representative agent is known and that it depends on the states of the other agents through the mean-field distribution. [^1] Our framework easily extends to the unknown reward case by learning the reward function similarly to the unknown dynamics and estimating its epistemic uncertainty (see [12]). Before every episode $n$, the representative agent selects a non-stationary policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})\in\Pi$ where individual policies [^2] We use the term policy profile to indicate non-stationary policies, while we refer to their components as individual policies. are of the form $\pi_{n,t}:\mathcal{S}\times\mathcal{P}(\mathcal{S})\rightarrow\mathcal{A}$ and $\Pi$ is the set of admissible policy profiles. The policy profile is then shared with all the agents that choose their actions according to $\bm{\pi}_{n}$ during episode $n$.
We consider a general family of transition dynamics $f:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathcal{S}$ defined as follows. Given the current mean-field distribution $\mu_{n,t}$, the representative agent’s state $s_{n,t}$ and its action $a_{n,t}$, the next representative agent’s state $s_{n,t+1}$ is given by:
@@
where $\varepsilon_{n,t}$ is the Gaussian noise with known variance. We assume that the transition dynamics are unknown and are to be inferred from collected trajectories across episodes.
Mean-field transition dynamics. State-to-state transition map in Equation 1 naturally extends to a mean-field transition dynamics induced by a policy profile $\bm{\pi}_{n}$ and dynamics $f$ in episode $n$ (see [43, Lemma 1]):
@@
For a given policy profile $\bm{\pi}_{n}$ and mean-field distribution $\mu$, the performance of the representative agent at step $t$ is measured via its expected future reward for the rest of the episode, i.e., $\mathbb{E}\left[\sum_{j=t}^{T-1}r(s_{n,j},\mu_{n,j},a_{n,j})\rvert\mu_{n,t}=\mu\right]$. Here, the expectation is over the randomness in the state transitions as well as over the sampling of the representative agent’s initial state, i.e., $s_{n,t}\sim\mu$.
3.1 Safe Mean-Field Reinforcement Learning
We define a safety function $h:\mathcal{P}(\mathcal{S})\to\mathbb{R}$ applied to the mean-field distribution as follows. Given some hard safety threshold $C\in\mathbb{R}$, we consider mean-field distributions $\mu$ as safe if they satisfy $h(\mu)\geq C$, or, equivalently, if
$h_{C}(\mu):=h(\mu)-C\geq 0.$ (3)
We denote as $\zeta=\{\mu\in\mathcal{P}(\mathcal{S}):h_{C}(\mu)\geq 0\}$, the safe mean-field distribution set for a safety constraint $h_{C}(\cdot)$. Hence, our focus is on the safety of the system as a whole rather than the safety of individual agents, as it becomes intractable handling individual agents’ states and interactions in the case of a large population. For the sake of exposition, we use a single constraint, however, our approach is directly extendable to multiple constraints.
For a given initial distribution $\mu_{0}$, we formally define the Safe Multi-agent Mean-field Reinforcement Learning problem as follows:
$\bm{\pi}^{*}$ $=\operatorname*{arg\,max}_{\bm{\pi}\in\Pi}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(s_{t},\mu_{t},a_{t})\Big{\rvert}\mu_{0}\Bigg{]}$ (4a)
[…]
$s_{t+1}$ $=f(s_{t},\mu_{t},a_{t})+\varepsilon_{t}$ (4c)

### v2

@@ Abstract.
Many applications, e.g., in shared mobility, require coordinating a large number of agents. Mean-field reinforcement learning addresses the resulting scalability challenge by optimizing the policy of a representative agent interacting with the infinite population of identical agents instead of considering individual pairwise interactions. In this paper, we address an important generalization where there exist global constraints on the distribution of agents (e.g., requiring capacity constraints or minimum coverage requirements to be met). We propose Safe-M3-UCRL, the first model-based mean-field reinforcement learning algorithm that attains safe policies even in the case of unknown transitions. As a key ingredient, it uses epistemic uncertainty in the transition model within a log-barrier approach to ensure pessimistic constraints satisfaction with high probability. Beyond the synthetic swarm motion benchmark, we showcase Safe-M3-UCRL on the vehicle repositioning problem faced by many shared mobility operators and evaluate its performance through simulations built on vehicle trajectory data from a service provider in Shenzhen. Our algorithm effectively meets the demand in critical areas while ensuring service accessibility in regions with low demand.
Key words and phrases:
Multi-agent reinforcement learning; Mean-field control; Global safety; Epistemic uncertainty; Probabilistic neural network ensemble; Shared mobility; Vehicle repositioning
@@
Figure 1. An illustration of vehicles’ spatial distribution (light-blue scatters), repositioning trips (blue arrows), and a trajectory of passenger trips (red arrows).
An illustration of vehicles’ spatial distribution (light-blue scatters), repositioning trips (blue arrows), and a trip trajectory as a sequence of passenger trips (red arrows)
Figure 1. An illustration of vehicles’ spatial distribution (light-blue scatters), repositioning trips (blue arrows), and a trajectory of passenger trips (red arrows).
1. Introduction
Multi-Agent Reinforcement Learning (MARL) is a rapidly growing field that seeks to understand and optimize the behavior of multiple agents interacting in a shared environment. MARL has a wide range of potential applications, including vehicle repositioning in shared mobility services (e.g., moving idle vehicles from low-demand to high-demand areas Lin et al. (2018)), swarm robotics (e.g., operating a swarm of drones Alon and Zhou (2020)), and smart grids (e.g., operating a network of sensors in electric system Roesch et al. (2020)). The interactions between agents in these complex systems introduce several challenges, including non-stationarity, scalability, competing learning goals, and varying information structure. Mean-Field Control (MFC) addresses the scalability and non-stationarity hurdles associated with MARL by exploiting the insight that many relevant MARL problems involve a large number of very similar agents working towards the same goal. Instead of focusing on the individual agents and their interactions, MFC considers an asymptotically large population of identical cooperative agents and models them as a distribution on the state space. This approach circumvents the problem’s dependency on the population size, enabling the consideration of large populations. The solutions obtained by MFC are often sufficient for the finite-agent equivalent problem Lacker (2017); Mondal et al. (2022); Wang et al. (2020); Chen et al. (2021) in spite of the introduced approximations. An example of such a system is a ride-hailing platform that serves on-demand trips with a fleet of vehicles. The platform needs to proactively reposition idle vehicles based on their current locations, the locations of the other vehicles in the fleet, and future demand patterns (see Figure 1) to maximize the number of fulfilled trips and minimize customer waiting times. Additionally, the platform may be obligated by external regulators to guarantee service accessibility across the entire service region. The problem quickly becomes intractable as the number of vehicles increases. A further difficulty lies in modeling the traffic flows. Due to numerous infrastructure, external, and driver behavioral factors, which are often region-specific, it is laborious and often difficult to determine transitions precisely Daganzo (1994); van der Gun et al. (2018); Bliemer and Raadsen (2020).
In this paper, we focus on learning the safe optimal policies for a large multi-agent system when the underlying transitions are unknown. In most real-world systems, the transitions must be learned from the data obtained from repeated interactions with the environment. We assume that the cost of obtaining data from the environment is high and seek to design a model-based solution that efficiently uses the collected data. Existing works consider solving the MFC problem via model-free or model-based methods without safety guarantees. However, the proposed Safe Model-Based Multi-Agent Mean-Field Upper-Confidence Reinforcement Learning (Safe-M3-UCRL) algorithm focuses on learning underlying transitions and deriving optimal policies for the mean-field setting while avoiding undesired or dangerous distributions of agents’ population.
@@ 1. Introduction
Contributions. Section 3 extends the MFC setting with safety constraints and defines a novel comprehensive notion of global population-based safety. To address safety-constrained environments featuring a large population of agents, in Section 4, we propose a model-based mean-field reinforcement learning algorithm called Safe-M3-UCRL. Our algorithm leverages epistemic uncertainty in the transition model, employing a log-barrier approach to guarantee pessimistic satisfaction of safety constraints and enables the derivation of safe policies. In Section 5, we conduct empirical testing on the synthetic swarm motion benchmark and real-world vehicle repositioning problem, a challenge commonly faced by shared mobility operators. Our results demonstrate that the policies derived using Safe-M3-UCRL successfully fulfill demand in critical demand hotspots while ensuring service accessibility in areas with lower demand.
2. Related Work
@@
Our notion of safety for the mean-field problem extends the frameworks of Mean-Field Games (MFG) and Mean-Field Control (MFC) Lasry and Lions (2006a, b); Huang et al. (2006); Huang et al. (2007). For a summary of the progress focusing on MFGs, see Laurière et al. (2022) and references therein. We focus on MFCs in this work, which assume cooperative agents in contrast to MFGs, which assume competition. Gast et al. (2012); Motte and Pham (2019); Gu et al. (2020, 2021a); Bäuerle (2021) address the problem of solving MFCs under known transitions, i.e., planning, while Wang et al. (2020); Carmona et al. (2019b); Angiuli et al. (2022, 2021); Carmona et al. (2019a); Wang et al. (2021); Carmona et al. (2021); Subramanian and Mahajan (2019) consider model-free Q-learning and Policy Gradient methods in various settings. Closest to our approach, Pásztor et al. (2023) introduces M3-UCRL, a model-based, on-policy algorithm, which is more sample efficient than other proposed approaches. Similarly to Jaksch et al. (2010); Chowdhury and Gopalan (2019); Curi et al. (2020) for model-based single-agent RL and Sessa et al. (2022) for model-based MARL, M3-UCRL uses the epistemic uncertainty in the transition model to design optimistic policies that efficiently balance exploration and exploitation in the environment and maximize sample efficiency. This is also the setting of our interest. However, safety is not considered in any of these methods.
In terms of safety, there are two main ways of handling it in RL; assigning significantly lower rewards to unsafe states Moldovan and Abbeel (2012) and providing additional knowledge to the agents Song et al. (2012) or using the notion of controllability to avoid unsafe policies explicitly Gehring and Precup (2013). Furthermore, the following approaches combine the two methods; Berkenkamp et al. (2017) uses Lyapunov functions to restrict the safe policy space, Cheng et al. (2019) projects unsafe policies to a safe set via a control barrier function, and Alshiekh et al. (2018) introduces shielding, i.e., correcting actions only if they lead to unsafe states. For comprehensive overviews on safe RL, we refer the reader to Garcıa and Fernández (2015); Gu et al. (2022). As an alternative, Usmanova et al. (2022) demonstrates that the general-purpose stochastic optimization methods can be used for constrained MDPs, i.e., safe RL formulations. Similar to our work, they use the log-barrier approach to turn constrained into unconstrained optimization. Nevertheless, the aforementioned works focus mainly on individual agents, while in large-scale multi-agent environments, maintaining individual safety becomes intractable, and the focus shifts towards global safety measures. For multi-agent problems, previous works focus on satisfying the individual constraints of the agents while learning in a multi-agent environment. For the cooperative problem, Gu et al. (2021b) proposes two model-free algorithms, MACPO and MAPPO-Lagrangian. MACPO is computationally expensive, while MAPPO-Lagrangian does not guarantee hard constraints for safety. Dec-PG solves the decentralized learning problem using a consensus network that shares weights between neighboring agents Lu et al. (2021). For the non-cooperative decentralized MARL problem with individual constraints, Sheebaelhamd et al. (2021) adds a safety layer to multi-agent DDPG Lowe et al. (2017) similar to single-agent Safe DDPG Dalal et al. (2018) for continuous state-action spaces. Aggregated and population-based constraints have been addressed in the following works. CMIX Liu et al. (2021) extends QMIX Rashid et al. (2020), which considers average and peak constraints defined over the whole population of agents in a centralized-learning decentralized-execution framework. Their formulation relies on the joint state and action spaces, making it infeasible for a large population of agents. ElSayed-Aly et al. (2021) introduces an additional shielding layer that corrects unsafe actions. Their centralized approach suffers from scalability issues, while the factorized shielding method monitors only a subset of the state or action space. For mixed cooperative-competitive settings, Zhu et al. (2020) uses the notion of returnability to define a safe, decentralized, multi-agent version of Q-learning that ensures individual and joint constraints. However, their approach requires an estimation of other agents’ policies, which does not scale well for large systems. Works considering constraints on the whole population fail to overcome the exponential nature of multi-agent problems or require domain knowledge to factorize the problems into subsets.
Closest to our setting, Mondal et al. (2022) introduces constraints to the MFC by defining a cost function and a threshold that the discounted sum of costs can not exceed. We propose a different formulation that restricts the set of feasible mean-field distributions at every step, therefore, addressing the scalability issue and allowing for more specific control over constraints and safe population distributions.
3. Problem Statement
@@
Formally, we consider the episodic setting, where episodes $n{}={}1,\ldots,N$ each have $t{}={}0,\ldots,T-1$ discrete steps and the terminal step $t=T$. The state space $\mathcal{S}\subseteq\mathbb{R}^{p}$ and action space $\mathcal{A}\subseteq\mathbb{R}^{q}$ are the same for every agent. We use $s_{n,t}^{(i)}\in\mathcal{S}$ and $a_{n,t}^{(i)}\in\mathcal{A}$, to denote the state and action of agent $i\in\{1,\ldots,m\}$ in episode $n$ at step $t$. For every $n$ and $t$, the mean-field distribution $\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ describes the global state with $m$ identical agents when $m\rightarrow{}+{}\infty$, i.e.,
$\mu_{n,t}(ds){}={}\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{n,t}^{(i)}\in ds),$
@@
We consider the MFC model to capture a collective behavior of a large number of collaborative agents operating in the shared stochastic environment. This model assumes the limiting regime of infinitely many agents and homogeneity. Namely, all agents are identical and indistinguishable, therefore, solving MFC amounts to finding an optimal policy for a single, so-called, representative agent. The representative agent interacts with the mean-field distribution of agents instead of focusing on individual interactions and optimizes a collective reward. Due to the homogeneity assumption, the representative agent’s policy is used to control all the agents in the environment.
We posit that the reward $r:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathbb{R}$ of the representative agent is known and that it depends on the states of the other agents through the mean-field distribution. [^1] Our framework easily extends to unknown reward by estimating its epistemic uncertainty and learning it similarly to learning unknown transitions (see Chowdhury and Gopalan (2019)). Before every episode $n$, the representative agent selects a non-stationary policy profile $\bm{\pi}_{n}{}={}(\pi_{n,0},\ldots,\pi_{n,T-1})\in\Pi$ where individual policies are of the form $\pi_{n,t}:\mathcal{S}\times\mathcal{P}(\mathcal{S})\rightarrow\mathcal{A}$ and $\Pi$ is the set of admissible policy profiles. The policy profile is then shared with all the agents that choose their actions according to $\bm{\pi}_{n}$ during episode $n$.
We consider a general family of deterministic transitions $f:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathcal{S}$. Given the current mean-field distribution $\mu_{n,t}$, the representative agent’s state $s_{n,t}$ and its action $a_{n,t}$, the next representative agent’s state $s_{n,t{}+{}1}$ is given by
@@
where $\varepsilon_{n,t}$ is a Gaussian noise with known variance. We assume that the transitions are unknown and are to be inferred from collected trajectories across episodes.
Mean-field transitions. State-to-state transition map in Equation 1 naturally extends to the mean-field transitions induced by a policy profile $\bm{\pi}_{n}$ and transitions $f$ in episode $n$ (see (Pásztor et al., 2023, Lemma 1))
@@
For a given policy profile $\bm{\pi}_{n}$ and mean-field distribution $\mu$, the performance of the representative agent at step $t$ is measured via its expected future reward for the rest of the episode, i.e.,
$\mathbb{E}\left[\sum_{j=t}^{T-1}r(z_{n,j}){}\rvert{}\mu_{n,t}{}={}\mu\right].$
Here, the expectation is over the randomness in the transitions and over the sampling of the initial state, i.e., $s_{n,t}\sim\mu$.
3.1. Safe Mean-Field Reinforcement Learning
We extend the MFC with global safety constraints, [^2] For the exposition, we use a single constraint, however, our approach is directly extendable to multiple constraints. i.e., the constraints imposed on the mean-field distributions. We consider safety functions $h:\mathcal{P}(\mathcal{S})\to\mathbb{R}$ over probability distributions. Given some hard safety threshold $C\in\mathbb{R}$, we consider a mean-field distribution $\mu$ as safe if it satisfies $h(\mu){}\geq{}C$, or, equivalently
$h_{C}(\mu){}:={}h(\mu)-C{}\geq{}0.$ (3)
We denote the set of safe mean-field distributions for a safety constraint $h_{C}(\cdot)$ as $\zeta{}={}\{\mu\in\mathcal{P}(\mathcal{S}):h_{C}(\mu){}\geq{}0\}$. Hence, our focus is on the safety of the system as a whole rather than the safety of individual agents, as it becomes intractable to handle individual agents’ states and interactions in the case of a large population.
For a given initial distribution $\mu_{0}$, we formally define the Safe-MFC [^3] We refer to formulations under known transitions as control problems, while we reserve the term reinforcement learning for formulations under unknown transitions. problem as follows
$\bm{\pi}^{*}{}={}\operatorname*{arg\,max}_{\bm{\pi}\in\Pi}{}$ $\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(z_{t})\Big{\rvert}\mu_{0}\Bigg{]}$ (4a)
subject $\text{subject to}\quad a_{t}$ ${}={}\pi_{t}(s_{t},\mu_{t})$ (4b)
$s_{t{}+{}1}$ ${}={}f(z_{t}){}+{}\varepsilon_{t}$ (4c)